Henry bought 32 CDs and DVDs for $1010. The CDs cost $25 each and the DVDs cost $40 each, How Many more CDs than DVDs did he buy.
Answer by ewatrrr(24785)

x the number of DVDs
40x + 25(32-x) = $1010
x =( 1010 - 800)/15 = 14, the number of DVDs. 18 CDs 32-14
...
And...checking
$560 + $450 = $1010